# Contamination Detection for VLMs Using Multi‑Modal Semantic Perturbations

### v1

@@ Abstract
Recent advances in Vision–Language Models (VLMs) have achieved state-of-the-art performance on numerous benchmark tasks. However, the use of internet-scale, often proprietary, pretraining corpora raises a critical concern for both practitioners and users: inflated performance due to test-set leakage. While prior works have proposed mitigation strategies such as decontamination of pretraining data and benchmark redesign for LLMs, the complementary direction of developing detection methods for contaminated VLMs remains underexplored. To address this gap, we deliberately contaminate open-source VLMs on popular benchmarks and show that existing detection approaches either fail outright or exhibit inconsistent behavior. We then propose a novel simple yet effective detection method based on multi-modal semantic perturbation, demonstrating that contaminated models fail to generalize under controlled perturbations. Finally, we validate our approach across multiple realistic contamination strategies, confirming its robustness and effectiveness. The code and perturbed dataset will be released publicly.
@@ 1 Introduction
Recent advances in Vision-Language Models (VLMs) have resulted in remarkable performance across a wide range of tasks, including visual reasoning (mmmu; mmbench; mmstar), real-world understanding (realworldqa), and complex mathematical problems (mathverse; mathvista). A typical VLM training pipeline involves pretraining a vision encoder and language backbone on internet-scale data, followed by a fine-tuning stage on high-quality multimodal instruction-tuning datasets. However, as these training corpora are often proprietary with their exact composition undisclosed, a critical concern has emerged: public benchmark data may have leaked into the training set, leading to inflated and misleading performance metrics.
@@
N-gram Accuracy (ngram-acc)
Shared Likelihood (provingtestset)
Guided Prompting (guided-prompting)
Multi-modal Leakage (mmstar) $\blacktriangle$
CircularEval (mmbench) $\blacktriangle$
Choice Confusion (yao-dcclb) $\blacktriangle$
BGR Shuffling (vlm-cleaneval)
Image Masking / Option Shuffling (vlm-systematic)
@@
Table 1: Analysis of existing detection methods on VLMs. We label the detection method with if it satisfies all of Requirement Requirement 1., Requirement 2. or Requirement 3. and with otherwise. $\blacktriangle$ indicates that the requirement is partially observed but not consistently with varying contamination settings. Most existing detection methods fail to meet the requirements and cannot accurately classify contaminated models. Our method, however, satisfies all requirements. Results for N-gram Accuracy, Shared Likelihood and Guided Prompting are in the Appendix B.
Test-set leakage presents a practical and significant challenge. For model users, it becomes difficult to disentangle genuine reasoning and generalization from mere memorization. For developers, exhaustively verifying the absence of test examples in massive pretraining corpora is prohibitively expensive  (qwentechnical). Although early work on large models proposed decontamination steps by removing n-gram overlaps (gpt3; abdin2024phi4technicalreport), many recent models do not report such procedures, leaving the extent of contamination largely unexamined.
To address this, several methods have been proposed to detect data contamination. One line of work focuses on verbatim memorization, testing whether a model can reconstruct exact benchmark questions with high confidence (ngram-acc; guided-prompting; provingtestset). Another line of work examines generalization failures, measuring if a model that succeeds on an original question fails on a simpler variant, which is interpreted as an evidence of memorization (gsm-symbolic; yao-dcclb; mathperturb).
@@
In this study, we conduct a systematic analysis of the data contamination problem, from which we derive grounded definitions of the core requirements – reliability, practicality, and consistency. Guided by these definitions, we systematically contaminate open-source VLMs under varying fine-tuning epochs, data composition, and training strategies (e.g., standard fine-tuning vs. LoRA (lora)). Our results show that existing detection methods struggle with the complexities of VLMs, frequently failing to meet the core requirements across diverse contamination scenarios (see Table 1).
@@
    We validate our method across multiple contamination settings, proving that it is reliable, practical, and consistent, satisfying all key requirements for a robust detection method.
@@ 2 Investigation Setup
This definition motivates our central assumption, grounded in prior work on model memorization (zhang2017; carlini; kandpal):
@@
In particular, we posit that as the degree of contamination grows, models will exhibit degraded performance on perturbed or out-of-distribution variants of benchmark items, even when the original questions are answered correctly. The effect need not scale linearly, since fine-tuning can disproportionately distort embedding spaces (choi-conta).
@@
Requirement 1. (Practicality). The method must operate without the knowledge of the leaked data or the model’s training corpus, relying only on black-box interactions. Requirement 2. (Reliability). The method must detect contaminated models across heterogeneous fine-tuning strategies (e.g., standard fine-tuning vs. LoRA). Requirement 3. (Consistency). The method’s detection signal should be positively correlated with degree of contamination (i.e., $n=\text{deg}_{\mathcal{D}}(\mathcal{M})$).
Together, these requirements define a principled framework for evaluating contamination detection. A method that satisfies all of them can practically flag contaminated models, remain agnostic to training specifics, and provide a signal proportional to the extent of contamination.
@@ 3 Preparation of Contaminated Models
To analyze contamination detection across diverse settings, we pair complementary model families and benchmarks. On the model side, we use LLaVA-v1.5-7B (llava1.5) – a LLaMA–adapter design – and Qwen2-VL-7B (qwen2), which integrates tighter multimodal alignment. Both are widely adopted open-source VLMs with publicly documented training details (e.g., released corpora or decontamination policies, although we do not rely on this for detection), making them suitable for controlled contamination experiments. Spanning these distinct training paradigms lets us probe robustness under heterogeneous contamination footprints.
For benchmarks, we use MMStar (mmstar) and RealWorldQA (realworldqa), which consist of questions that strictly require visual and textual evidence. MMStar aggregates and filters prior multi-modal benchmarks by removing questions that are leaked or do not have visual dependence, while RealWorldQA enforces visual dependence by design. This prevents linguistic shortcuts and ensures our perturbation-based detector evaluates genuine multi-modal reasoning.
@@
We compare standard fine-tuning and LoRA (lora). For standard fine-tuning, we follow three common variants: fine-tune the LLM and adapter as in LLaVA-v1.5-7B; fine-tune only the LLM as in Qwen2-VL-7B; or unfreeze all parameters as in InternVL (internvl). This diversity tests robustness across parameter-efficient and full fine-tuning regimes.
@@
We save checkpoints at epochs 1–3 (i.e. $\deg_{\mathcal{D}}(\mathcal{M})\in\{1,2,3\}$), since VLMs are typically fine-tuned for at most three epochs – and often just one. Contamination footprints of varying strength enables a graded analysis of detection sensitivity (llava1.5; internvl2.5).
@@
We use model defaults, adjusting only the learning rate to ensure inflated performance. For LLaVA-v1.5-7B, we follow the official repo; for Qwen2-VL-7B, we use LLaMA-Factory (llamafactory) to train the models. (Full settings are in the Appendix A.1.)
@@ 4 Multi-Modal Semantic Perturbation
To generate semantically perturbed questions, we combine an LLM with a diffusion-based generative model. In our main experiments, we use GPT-4o (gpt4o) and Flux (flux2024) + ControlNet (Zhang2023ConditionalControl), but later show that our framework is model-agnostic (Section  7).
The pipeline of our framework is as follows. First, we randomly change the answer of the original question to a different option, preventing contaminated models from getting away with memorized responses. Next, GPT-4o generates a dense caption of the image, conditioned both on the original question and the newly chosen answer choice. We find this explicit conditioning essential: it ensures the caption highlights the salient visual features that drive the perturbation process, allowing the generation of questions of similar or lower difficulty (Section  6). Flux ControlNet then uses this caption, together with Canny edge maps (canny), to guide the diffusion process. ControlNet preserves the global structure of the image while introducing new elements that minimally alter semantics, yielding an updated image with the randomly sampled different correct answer (Figure 2). We note that, due to limitations in rendering text or complex geometries, especially at low resolution, some generated images do not fully correspond to the new answer. To mitigate this, we filter generated pairs using a single criterion: perturbed questions must be answerable unambiguously. [^1] To further clarify, generation quality itself is not considered in filtering. This ensures that the evaluation set focuses solely on reasoning rather than visual fidelity. While our main results report human-filtered outcomes, we show in Table 7 that this step can be automated. We stress that filtering is necessitated by current generative model limitations, not by our detection principle.
To detect contamination, i.e., memorization of an image-text pair, we compare model performance on the original vs. perturbed input. If a model fails to generalize (i.e., it gets the original input correct but the perturbed one incorrect), we flag contamination. A clean model with genuine reasoning capabilities should perform correctly in both settings, given that they have comparable difficulty. Critically, our approach enables robust contamination detection without requiring any ground-truth knowledge of the leaked data.
@@ 5 Comparative Evaluation of Contamination Detection
First, we see that clean models perform better than contaminated models, confirming that the perturbed questions are indeed of equal or lower difficulty. In contrast, all contaminated models perform worse, enabling our method to reliably detect via a simple check for performance drops.
@@
Table 2: Performance of LLaVA-v1.5-7B (top) and Qwen2-VL-7B (bottom) on the MMStar dataset (Corresponding RealWorldQA results are in the Appendix 11). We compare to “Multi‐modal Leakage” mmstar, CircularEval (mmbench), and “Choice Confusion” yao-dcclb. Clean models perform better on our perturbed dataset – confirming that the perturbed questions are indeed of equal or lower difficulty. In contrast, all contaminated models perform worse, enabling reliable detection by our method via a simple check for performance drops. "_P" denotes the semantically perturbed version; "to" denotes text-only performance; "_C" denotes evaluation using circular options; "_G" denotes evaluation using choice confusion; $\Delta$ denotes the difference in performance, with positive values indicating gains. "Success?" indicates whether the method detected contamination. "Require clean model?" indicates whether the method requires access to a clean model as a baseline. If a clean model is required as reference, the method cannot be used to detect the reference models themselves, so the entry is marked as “–”. Full results of multi-modal semantic perturbation is listed in Appendix E.1.
@@
Multi-modal Leakage (mmstar). With multi-modal leakage, contamination is flagged by measuring text‐only performance on benchmarks that require visual input. After fine‐tuning, any boost in text‐only performance indicates memorization of leaked question–answer pairs rather than genuine multi-modal reasoning. Multi-modal leakage requires clean models by design and hence does not satisfy Practicality (Req. Requirement 1.). And from Table 2 we observe that Reliability (Req. Requirement 2.) is not satisfied since it fails to detect LLaVA-v1.5-7B  trained for 3 epochs with standard fine-tuning. While its gain in performance positively correlates with the degree of contamination in other cases, this trend breaks across benchmarks and training strategies such as for Qwen2-VL-7B trained for 3 epochs with LoRA and LLaVA-v1.5-7B  trained for 3 epochs with standard fine-tuning, thus only partially satisfying Consistency (Req. Requirement 3.).
@@
CircularEval (mmbench). Selection bias in multiple-choice questions is mitigated by rotating answer options $n$ times and requiring the model to be correct on all rotations. This stricter criterion lowers absolute accuracy compared to standard evaluation. For contamination detection, however, CircularEval’s effectiveness is limited. Practicality (Req. Requirement 1.) fails, as CircularEval lacks a clear threshold-independent detection mechanism. Reliability (Req. Requirement 2.) is undermined by inconsistent contamination signals, as it fails to detect LLaVA-v1.5-7B  trained with LoRA for 2 and 3 epochs, and Qwen2-VL-7B  trained with standard fine-tuning for 2 epochs. And despite some positive trends, Consistency (Req. Requirement 3.) breaks across models and training setups, as shown with LLaVA-v1.5-7B  trained with LoRA for 2, 3 epochs and Qwen2-VL-7B  trained with standard fine-tuning for 2,3 epochs, limiting its diagnostic value.
@@
Choice confusion (yao-dcclb). Generalization is tested by constructing an easier benchmark variant: false options are replaced with correct answers drawn from unrelated questions. Clean models should leverage this simplification to improve, whereas contaminated models—bound by memorized original answers—might not. We apply this method to measure the model’s performance drop ($\Delta$) between the original and generalized versions. Clean models gain substantially on generalized benchmarks – up to +34.04 on MMStar and +21.30 on RQA – confirming their generalization ability. By contrast, contaminated variants show much smaller gains or even losses. This failure to benefit from semantically irrelevant but easier choices reflects classic memorization‐based contamination. As a detection method, choice confusion meets Practicality (Req. Requirement 1.) since the method detects models that perform worse on the generalized benchmark, but its sensitivity varies across fine-tuning regimes. Choice confusion fails to detect LLaVA-v1.5-7B  regardless of its training strategy or the number of epochs, and for LLaVA-v1.5-7B  trained with LoRA, the model shows improved performance as the number of training epochs increases. Hence the method fails to satisfy Reliability (Req. Requirement 2.) while Consistency (Req. Requirement 3.) is partially observed in other training strategies.
@@
Importantly, unlike other approaches, our method requires no dataset-specific thresholds or prior knowledge of leaked data, satisfying Practicality (Req. Requirement 1.). Moreover, the performance drop scales with contamination degree, satisfying Consistency (Req. Requirement 3.), and persists across all training regimes—standard fine-tuning, LoRA, and full parameter unfreezing, satisfying Reliability (Req. Requirement 2.).

### v2

@@ Abstract
Recent advances in Vision–Language Models (VLMs) have achieved state-of-the-art performance on numerous benchmark tasks. However, the use of internet-scale, often proprietary, pretraining corpora raises a critical concern for both practitioners and users: inflated performance due to test-set leakage. While prior works have proposed mitigation strategies such as decontamination of pretraining data and benchmark redesign for LLMs, the complementary direction of developing detection methods for contaminated VLMs remains underexplored. To address this gap, we deliberately contaminate open-source VLMs on popular benchmarks and show that existing detection approaches either fail outright or exhibit inconsistent behavior. We then propose a novel simple yet effective detection method based on multi-modal semantic perturbation, demonstrating that contaminated models fail to generalize under controlled perturbations. Finally, we validate our approach across multiple realistic contamination strategies, confirming its robustness and effectiveness. The code and perturbed dataset are released here: https://github.com/jadenpark0/mm-perturb.
@@ 1 Introduction
Recent advances in Vision-Language Models (VLMs) have resulted in remarkable performance across a wide range of tasks, including visual reasoning (Yue et al., 2024; Liu et al., 2024b; Chen et al., 2024a), real-world understanding (xAI, 2024), and complex mathematical problems (Zhang et al., 2024b; Lu et al., 2024b). A typical VLM training pipeline involves pretraining a vision encoder and language backbone on internet-scale data, followed by a fine-tuning stage on high-quality multimodal instruction-tuning datasets. However, as these training corpora are often proprietary with their exact composition undisclosed, a critical concern has emerged: public benchmark data may have leaked into the training set, leading to inflated and misleading performance metrics.
@@
N-gram Accuracy (Xu et al., 2024)
Shared Likelihood (Oren et al., 2023)
Guided Prompting (Golchin and Surdeanu, 2024)
Multi-modal Leakage (Chen et al., 2024a) $\blacktriangle$
CircularEval (Liu et al., 2024b) $\blacktriangle$
Choice Confusion (Yao et al., 2024) $\blacktriangle$
BGR Shuffling (Lu et al., 2024a)
Image Masking / Option Shuffling (Song et al., 2025)
@@
Table 1: Analysis of existing detection methods on VLMs. We label the detection method with if it satisfies all of Requirement Requirement 1., Requirement 2. or Requirement 3. and with otherwise. $\blacktriangle$ indicates that the requirement is partially observed but not consistently with varying contamination settings. Most existing detection methods fail to meet the requirements and cannot accurately classify contaminated models. Our method, however, satisfies all requirements. Results for N-gram Accuracy, Shared Likelihood and Guided Prompting are in the Appendix C.
Test-set leakage presents a practical and significant challenge. For model users, it becomes difficult to disentangle genuine reasoning and generalization from mere memorization. For developers, exhaustively verifying the absence of test examples in massive pretraining corpora is prohibitively expensive (Bai et al., 2023). Although early works on large models proposed decontamination steps by removing n-gram overlaps (Brown et al., 2020; Abdin et al., 2024a), many recent models do not report such procedures, leaving the extent of contamination largely unexamined.
To address this, several methods have been proposed to detect data contamination. One line of work focuses on verbatim memorization, testing whether a model can reconstruct exact benchmark questions with high confidence (Xu et al., 2024; Golchin and Surdeanu, 2024; Oren et al., 2023). Another line of work examines generalization failures, measuring if a model that succeeds on an original question fails on variants with similar or easier difficulty, which is interpreted as an evidence of memorization (Mirzadeh et al., 2024; Yao et al., 2024; Huang et al., 2025).
@@
In this study, we conduct a systematic analysis of the data contamination problem, from which we derive grounded definitions of the core requirements – reliability, practicality, and consistency. Guided by these definitions, we systematically contaminate open-source VLMs under varying fine-tuning epochs, data composition, and training strategies (e.g., standard fine-tuning vs. LoRA (Hu et al., 2021)). Our results show that existing detection methods struggle with the complexities of VLMs, frequently failing to meet the core requirements across diverse contamination scenarios (see Table 1).
@@
    We validate our method across multiple contamination settings, proving that it is reliable, practical and consistent, satisfying all key requirements for a robust detection method.
@@ 2 Investigation Setup
This definition motivates our central assumption, grounded in prior work on model memorization (Zhang et al., 2017; Carlini et al., 2021; Kandpal et al., 2023):
@@
In particular, we posit that as the degree of contamination grows, models will exhibit degraded performance on perturbed or out-of-distribution variants of benchmark items, even when the original questions are answered correctly. The effect need not scale linearly, since fine-tuning can disproportionately distort embedding spaces (Choi et al., 2025).
@@
Requirement 1. (Practicality). The method must operate without assuming any prior knowledge of clean model (e.g. the model’s training corpus), relying only on black-box interactions. Requirement 2. (Reliability). The method must detect contaminated models across heterogeneous fine-tuning strategies (e.g., standard fine-tuning vs. LoRA). Requirement 3. (Consistency). The method’s detection signal should be positively correlated with degree of contamination $n=\text{deg}_{\mathcal{D}}(\mathcal{M})$.
Together, these requirements define a principled framework for evaluating contamination detection. A method that satisfies all of them can practically flag contaminated models without the knowledge of clean model’s behavior, remain agnostic to training specifics, and provide a signal proportional to the extent of contamination.
@@ 3 Preparation of Contaminated Models
To analyze contamination detection across diverse settings, we pair complementary model families and benchmarks. On the model side, we use LLaVA-v1.5-7B (Liu et al., 2024a) – a LLaMA–adapter design – and Qwen2-VL-7B (Wang et al., 2024), which integrates tighter multimodal alignment. Both are widely adopted open-source VLMs with publicly documented training details (e.g., released corpora or decontamination policies, though we do not rely on this for detection), making them suitable for controlled contamination experiments.
For benchmarks, we use MMStar (Chen et al., 2024a) and RealWorldQA (xAI, 2024), which consist of questions that strictly require visual and textual evidence. MMStar aggregates and filters prior multi-modal benchmarks by removing questions that are leaked or do not have visual dependence, while RealWorldQA enforces visual dependence by design. This prevents linguistic shortcuts and ensures our perturbation-based detector evaluates genuine multi-modal reasoning.
@@
We compare standard fine-tuning and LoRA (Hu et al., 2021). For standard fine-tuning, we follow three common variants: fine-tune the LLM and adapter as in LLaVA-v1.5-7B; fine-tune only the LLM as in Qwen2-VL-7B; or unfreeze all parameters as in InternVL (Chen et al., 2024b). This diversity tests robustness across parameter-efficient and full fine-tuning regimes.
@@
We save checkpoints at epochs 1–3 (i.e. $\deg_{\mathcal{D}}(\mathcal{M})\in\{1,2,3\}$), since VLMs are typically fine-tuned for at most three epochs – and often just one. Contamination footprints of varying strength enables a graded analysis of detection sensitivity (Liu et al., 2024a; Chen et al., 2025).
@@
We use model defaults, adjusting only the learning rate to ensure inflated performance. For LLaVA-v1.5-7B, we follow the official repo; for Qwen2-VL-7B, we use LLaMA-Factory (Zheng et al., 2024) to train the models. (Full settings are in the Appendix A.1.)
@@ 4 Multi-Modal Semantic Perturbation
To generate semantically perturbed questions, we combine an LLM with a diffusion-based generative model. In our main experiments, we use GPT-4o (OpenAI, 2024) and Flux (Labs, 2024) + ControlNet (Zhang et al., 2023), but later show that our framework is model-agnostic (Section  7).
The pipeline is as follows. First, we randomly change the answer of the original question to a different option, preventing contaminated models from getting away with memorized responses. Next, GPT-4o generates a dense caption of the image, conditioned both on the original question and the newly chosen answer choice. We find this explicit conditioning essential: it ensures the caption highlights the salient visual features that drive the perturbation process, allowing the generation of questions of similar or lower difficulty (Section  6). Flux ControlNet then uses this caption, together with Canny edge maps (Canny, 1986), to guide the diffusion process. ControlNet preserves the global structure of the image while introducing new elements that minimally alter semantics, yielding an updated image with the randomly sampled different correct answer (Figure 2). We note that, due to limitations in rendering text or complex geometries, especially at low resolution, some generated images do not fully correspond to the new answer. To mitigate this, we filter generated pairs using a single criterion: perturbed questions must be answerable unambiguously. [^1] Generation quality itself is not considered. This ensures that the evaluation set focuses solely on reasoning rather than visual fidelity. While our main results report human-filtered outcomes, we show in Table 9 that this step can be automated. We stress that filtering is necessitated by current generative model limitations, not by our detection principle. For the main results, we apply manual filtering to demonstrate the upper-bound performance of our approach. However, as shown in Table 9, this manual step can be replaced with automated filtering with a strong reasoning model.
To detect whether a model is contaminated on a dataset-level, i.e. has memorized the training image-text pairs, we compare its aggregate performance (i.e. accuracy) on the original vs. perturbed benchmarks. If the model fails to generalize (i.e. it gets the original input correct but the perturbed one incorrect, leading to a lower performance on the perturbed benchmark), we flag contamination [^2] We make a note that our approach can also be used to detect contamination on a sample-level by comparing the model’s behavior on the perturbed sample instead of the aggregate performance.. This is because a clean model with genuine reasoning capabilities should perform correctly in both settings, given that they have comparable difficulty. Critically, our approach enables robust contamination detection without requiring any ground-truth knowledge of the clean model’s behavior.
@@ 5 Comparative Evaluation of Contamination Detection
First, we see that clean models perform better than contaminated models, confirming that the perturbed questions are indeed of equal or lower difficulty, and that we are testing against perturbations that the clean models are robust against. In contrast, all contaminated models perform worse, enabling our method to reliably detect via a simple check for performance drops.
@@
Table 2: Performance of LLaVA-v1.5-7B (top) and Qwen2-VL-7B (bottom) on the MMStar dataset (Corresponding RealWorldQA results are in the Appendix 13). We compare to “Multi‐modal Leakage” Chen et al. (2024a), CircularEval (Liu et al., 2024b), and “Choice Confusion” Yao et al. (2024). Clean models perform better on our perturbed dataset – confirming that the perturbed questions are indeed of equal or lower difficulty. In contrast, all contaminated models perform worse, enabling reliable detection by our method via a simple check for performance drops. "_P" denotes the semantically perturbed version; "to" denotes text-only performance; "_C" denotes evaluation using circular options; "_G" denotes evaluation using choice confusion; $\Delta$ denotes the difference in performance, with positive values indicating gains. "Success?" indicates whether the method detected contamination. "Require clean model?" indicates whether the method requires access to a clean model as a baseline. If a clean model is required as reference, the method cannot be used to detect the reference models themselves, so the entry is marked as “–”. Full results of multi-modal semantic perturbation is listed in Appendix F.
@@
Multi-modal leakage (Chen et al., 2024a). With multi-modal leakage, contamination is flagged by measuring text‐only performance on benchmarks that require visual input. Any boost in text‐only performance after fine-tuning indicates memorization of leaked problems rather than genuine multi-modal reasoning. Multi-modal leakage requires clean models by design and hence does not satisfy Practicality (Req. Requirement 1.). From Table 2 we observe that Reliability (Req. Requirement 2.) is not satisfied since it fails to detect LLaVA-v1.5-7B trained for 3 epochs with standard fine-tuning. While the performance gain positively correlates with the degree of contamination in other cases, this breaks across benchmarks and training strategies such as for Qwen2-VL-7B trained for 3 epochs with LoRA and LLaVA-v1.5-7B  trained for 3 epochs with standard fine-tuning, only partially satisfying Consistency (Req. Requirement 3.).
@@
CircularEval (Liu et al., 2024b). Selection bias in multiple-choice questions is mitigated by rotating answer options $n$ times and requiring the model to be correct on all rotations. This stricter criterion lowers absolute accuracy compared to standard evaluation. For contamination detection, however, CircularEval’s effectiveness is limited. Practicality (Req. Requirement 1.) fails, as CircularEval lacks a clear threshold-independent detection mechanism. Reliability (Req. Requirement 2.) is undermined by inconsistent contamination signals, as it fails to detect LLaVA-v1.5-7B  trained with LoRA for 2 and 3 epochs, and Qwen2-VL-7B  trained with standard fine-tuning for 2 epochs. And despite some positive trends, Consistency (Req. Requirement 3.) breaks across models and training setups, as shown with LLaVA-v1.5-7B  trained with LoRA for 2, 3 epochs and Qwen2-VL-7B  trained with standard fine-tuning for 2,3 epochs, limiting its diagnostic value.
@@
Choice confusion (Yao et al., 2024). Generalization is tested by constructing an easier benchmark variant: false options are replaced with correct answers drawn from unrelated questions. Clean models should leverage this simplification to improve, whereas contaminated models – bound by memorized original answers – might not. We apply this method to measure the model’s performance drop ($\Delta$) between the original and generalized versions. Clean models gain substantially on generalized benchmarks – up to +34.04 on MMStar and +21.30 on RQA – confirming their generalization ability. By contrast, contaminated variants show much smaller gains or even losses. This failure to benefit from semantically irrelevant but easier choices reflects classic memorization‐based contamination. As a detection method, choice confusion meets Practicality (Req. Requirement 1.) since the method detects models that perform worse on the generalized benchmark, but its sensitivity varies across fine-tuning regimes. Choice confusion fails to detect LLaVA-v1.5-7B  regardless of its training strategy or the number of epochs, and for LLaVA-v1.5-7B  trained with LoRA, the model shows improved performance as the number of training epochs increases. Hence the method fails to satisfy Reliability (Req. Requirement 2.) while Consistency (Req. Requirement 3.) is partially observed in other training strategies.
@@
Importantly, unlike other approaches, our method requires no dataset-specific thresholds or prior knowledge of leaked data, satisfying Practicality (Req. Requirement 1.). Moreover, the performance drop scales with contamination degree, satisfying Consistency (Req. Requirement 3.), and persists across all training regimes – standard fine-tuning, LoRA, and full parameter unfreezing, satisfying Reliability (Req. Requirement 2.).